OpenStudy (anonymous):

Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side? Use the appropriate formula to find the solution and round to the nearest tenth.

OpenStudy (campbell_st):

this pythagoras' theorem call the other short side a then \[a^2 = 15^2 - 8^2\] once you have that answer... take the square root... to get the necessary measurement.

OpenStudy (anonymous):

so solve a^2=15^2-8^2?